OpenStudy (anonymous):

Solving the equation 3^3x-2 = 3^5-2x for x, you get x = A)7/5 B)log3 7/5 c)7/3 d)log3 3/5 e)7

OpenStudy (anonymous):

that's : \[3^{3x-2} = 3^{5-2x}\]

OpenStudy (anonymous):

\[\LARGE 3^{3x-2}=3^{5-2x} \quad \iff 3x-2=5-2x\] so when you have the same bases just take exponents... 3x-2=5-2x 3x+2x=5+2 5x=7 can you do it now ? :)

OpenStudy (anonymous):

7/5 ....thanks. I actually did it the long way and i didn't want to submit just to make sure.

OpenStudy (anonymous):

:) ... My pleasure.